Claire is twice as old as her brother Derek. Ten years ago, the sum of their ages was 46. How old is each now?
Answer by vasumathi(46)

Let the age of Claire = c and the age of her brother = b
c = 2b ------------>(1)(since Claire is twice as old as her brother)
Ten years ago the age of Claire = c-10 and the age of her brother = b-10.And the
sum of these ages = 46.This implies ....
c-10 + b-10 = 46
c+b-20 = 46
c+b = 46+20 (adding 20 on both sides)
2b + b = 66 (from the (1) above)
3b = 66
b = 66/3 = 22
c = 2b = 2*22 = 44
So the age of Claire = 44 years
and the age of her brother = 22 years.